The point is the third vertex of a triangle with area 28 whose other two vertices are (-4, 2) and (4,2). What is the point(s)
Answer by ikleyn(52747)

.
The two given points are endpoints of the horizontal segment y = 2.


They define the side of a triangle of the length 4 - (-4) = 4 + 4 = 8 units.


Since the area of a triangle is 28 square units, its height is 7 units.


Therefore, the set of points under the questions are TWO horizontal straight lines parallel to the given line y = 2 
and remoted from this line at the distance of 7 units.


So, these lines are  y = 9  and  y = -5.

Solved, answered, explained and completed.